OpenStudy (anonymous):

5h-9=-16+6h A. 4 B. -7 c. 7 d. 10

HanAkoSolo (jamierox4ev3r):

first in these types of equation problems, you would subtract the like terms on both sides… can you name the like terms? @Aly.mac101 ?

OpenStudy (anonymous):

h=7 16-9=6h-5h

HanAkoSolo (jamierox4ev3r):

you might want to show a process first :|

OpenStudy (anonymous):

I think the answer is -7

HanAkoSolo (jamierox4ev3r):

@sbta2

HanAkoSolo (jamierox4ev3r):

@Aly.mac101 do you understand how that answer was obtained?

HanAkoSolo (jamierox4ev3r):

5h-9=-16+6h -5h -5h ------------- -9=-16+h +16 +16 ------- h=7

HanAkoSolo (jamierox4ev3r):

if you would like to check your answer, plug it in to the original equation (in place of "h")